OpenStudy (anonymous):

A doctor can prescribe one of two drugs to treat a certain disease. Drug A will cure 75% of the cases but has unwanted side effects in 40% of the cases. Drug B will cure 50% of the cases and has unwanted side effects in 30% of cases. 60 patients are given drug A and 80 patients are given drug B. How many patients would you expect to experience unwanted side effects? 30 48 49 85 88

OpenStudy (anonymous):

what is 40% of 60?

OpenStudy (anonymous):

24?

OpenStudy (anonymous):

yes \(24!\) what is 30% of 80?

OpenStudy (anonymous):

24? ha

OpenStudy (acxbox22):

24+24=?

OpenStudy (anonymous):

right and \(24+24=48\) i hope

OpenStudy (acxbox22):

i do too

OpenStudy (anonymous):

so 48%? thank you!!

OpenStudy (anonymous):

could have written \[4\times 6+3\times 8=48\] as well

OpenStudy (acxbox22):

not percent

OpenStudy (anonymous):

oooooh NO! not precent 48 people

OpenStudy (anonymous):

oh yeah whoops sorry

OpenStudy (anonymous):

yeah thank you sorry! :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

There are two prominent, nationally circulated newspapers, The Wall Street Journal and USA Today. Among all libraries across the country, 45% subscribe to The Wall Street Journal and 40% subscribe to USA Today, while only 25% subscribe to both The Wall Street Journal and USA Today. If a library is selected at random, what is the probability that it subscribes to The Wall Street Journal but not USA Today? 10% 15% 20% 25% 60%

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

In general, how can the volume of any prism or cylinder be found? a) Multiply length by width by height. b) Multiply length of the base by height of the figure. c) length by itself three times. d) Multiply area of the base by height of the figure.

OpenStudy (anonymous):

ill give a medal